Find the length of a rectangle of 131.2 square centimeters if the widthe is 8 centimeters.
Found 2 solutions by josmiceli, checkley71:
Answer by josmiceli(19441)

131.2 cm2 / 8 cm = 16.4 cm

Answer by checkley71(8403)

l*w=area
8*L=131.2
L=131.2/8
L=16.4 IS THE LENGTH.
PROOF
16.4*8=131.2
131.2=131.2